# Fast-Poly: A Fast Polyhedral Framework For 3D Multi-Object Tracking

### v1

@@ III-D Parallelization
Although the TBD framework is inherently serial, certain modules within it can be parallelized. Consequently, in Fast-Poly, following the tracking in the previous frame, the prediction in the estimation module and the preprocessing module are synchronized in time based on the multi-processing technology. This process is provided as a pseudo-code in Algorithm 1. Specifically, before the online tracking, two parallel processes $P_{1},P_{2}$ are initiated. This initialization is also accompanied by a shared flag $Flag$ and variable $V$ to synchronize status and transmit data between processes. At each frame $t$, $P_{1}$ filters raw detections while $P_{2}$ concurrently predicts the previous alive trajectory and implements online tracking. $P_{1}$ and the prediction part of $P_{2}$ enter a sleep state after completing their respective tasks, remaining dormant until they are simultaneously awakened in the subsequent frame following the completion of current frame tracking. Our approach adheres to online tracking principles, distinguishing itself fundamentally from [4], which employs multi-processing for synchronous multi-scene tracking.
@@ IV-A Dataset and Implementation Details
Metrics. We present four primary indicators: MOTA [23], AMOTA [10], HOTA [24], and FPS (Frame Per Second), supplemented by three secondary indicators: IDS (ID Switch), FP, and FN. The latency of each experiment is the average time cost under multiple evaluations.
TABLE I: A comparison between our proposed method with other advanced methods on the nuScenes test set. This leaderboard is available at nuScenes official benchmark. $\ddagger$ means the GPU device. The reported runtimes of all methods exclude the detection time. Fast-Poly and Poly-MOT [1] rely entirely on the detector input, as they do not utilize any visual or deep features.
Method Device Detector Input AMOTAnormal-$\uparrow$ MOTAnormal-$\uparrow$ FPSnormal-$\uparrow$ IDSnormal-$\downarrow$ FNnormal-$\downarrow$ FPnormal-$\downarrow$
@@
TABLE III: A comparison of the real-time performance on different devices on nuScenes val set with CenterPoint.
Device Level FPSnormal-$\uparrow$ AMOTAnormal-$\uparrow$ MOTAnormal-$\uparrow$
@@
Split Method Detector Input MOTAnormal-$\uparrow$
@@
TABLE V: A comparison between our proposed method with other advanced methods on the nuScenes val set. $\ddagger$ means the GPU device. All metrics in the competition papers are reported. Methods in lines 1-7 and lines 8-9 use CenterPoint [20] and LargekKernel3D-L [28], respectively.
Method Device AMOTAnormal-$\uparrow$ MOTAnormal-$\uparrow$ FPSnormal-$\uparrow$ IDSnormal-$\downarrow$
@@
TABLE VI: The ablation studies of each module on the nuScenes val. AG means A-GIoU. VM means Voxel Mask. LW means LightWeight Filter. CL means Confidence-Count Mixed Lifecycle. MP means Multi-Processing.
AG VM LW CL MP AMOTAnormal-$\uparrow$ MOTAnormal-$\uparrow$ FPSnormal-$\uparrow$ IDSnormal-$\downarrow$ FPnormal-$\downarrow$ FNnormal-$\downarrow$
Baseline [1] 73.1 61.9 5.6 (188.7ms) 232 13051 17593
Exp1 73.1 62.4 19.9 (50.3ms) 342 13955 16794
Exp2 73.3 63.0 21.3 (46.9ms) 385 12732 17909
Exp3 73.3 62.9 21.7 (46.1ms) 427 14124 16523
Exp4 73.7 63.2 26.5 (37.7ms) 414 14713 15900
Exp5 73.7 63.2 28.9 (34.6ms) 414 14713 15900
TABLE VII: A comparison on distinct lifecycle modules on nuScenes val set. Average means using the online average score to delete. Latest means using the latest score to delete. Max-age means using the continuous mismatch time to delete. Other settings are under the best performance.
Strategy AMOTAnormal-$\uparrow$ MOTAnormal-$\uparrow$ FPSnormal-$\uparrow$ FNnormal-$\downarrow$
Count & Max-age 73.3 62.9 23.0 16523
Confidence [15] & Latest 70.6 63.2 45.8 19192
Confidence [15] & Average 73.3 63.1 28.3 16826
Confidence (Ours) & Average 73.7 63.2 28.9 15900
IV-B Run-time Discussion
The robustness of device migration. As demonstrated in Table III, we evaluated Fast-Poly across three CPU tiers. Since our method is learning-free, tracking accuracy remains unaffected by device migration. On the personal-level CPU, Fast-Poly exhibits the best real-time performance. Besides, considering the challenging outdoor sequences of nuScenes, the inference speed (9 FPS) on Jetson Orin AGX enables its deployment in practical robotic systems.
The parallel efficiency of Fast-Poly. As depicted in Fig. 1, as the number of objects increases significantly, the average frame timing of Fast-Poly exhibits minimal increments ($\leq$30ms). This observation highlights the outstanding computational efficiency and feasibility of our method in autonomous driving scenarios.
@@ IV-C Comparative Evaluation
On the val set, we employ CenterPoint [20] as the detector for a fair comparison. As shown in Table V, Fast-Poly outperforms most learning-based/free methods by a significant margin in terms of both tracking accuracy (73.7% AMOTA, 64.2% MOTA) and latency (29 FPS). We surpass the baseline [1] with a notable improvement of +1.3% in MOTA and +0.6% in AMOTA accuracy, while 5$\times$ faster under identical settings. With a powerful detector LargeKernel3D [28], Fast-Poly exhibits stronger tracking performance.
Waymo. As shown in Table  IV, with the same detector, we exceed CasTrack [13] with a +1% improvement in MOTA on both test and val sets, showcasing superior tracking performance. Besides, without using the state-of-the-art detector, Fast-Poly ranks 4th on the Waymo tracking leaderboard among all LiDAR-only online methods with 63.6% MOTA to date. Despite the scarcity of methods showcasing real-time performance on Waymo, we believe that our inference speed (35.5 FPS) is competitive.
@@ IV-D Ablation Studies
The effect of the Proposed Metric A-GIoU. As demonstrated in Tables II and VI, $A\text{-}gIoU$ results in a significant reduction in latency (-138.4ms), particularly notable during the overlap (-24.7ms) and convex hull (-83ms) calculation phases. Additionally, an intriguing observation arises from Table VI, where $A\text{-}gIoU$ maintains AMOTA (73.1%) and also enhances MOTA (+0.5%). This enhancement stems from the alignment facilitating association to recall more observations (-799 FN).
The effect of the Voxel Mask. As illustrated in Table VI, the increase in accuracy (+0.2% AMOTA, +0.6% MOTA) and decrease in latency (-3.4ms) verify that the voxel mask can avoid invalid cost calculations. An important insight is that the voxel mask avoids excessively distant correlations while also limiting the mismatches (-1223 FP).
The effect of the Lightweight Filter. As indicated in Table VI, combing the lightweight filter results in enhancements in real-time performance (-0.8ms). It is worth mentioning that the lightweight filter also has a slight negative impact on accuracy (-0.1% MOTA), due to the increase in FP and IDS.
The effect of the Confidence-Count Mixed Lifecycle. The notable reduction in FN (-623) and the improvement in AMOTA (+0.4%) showcased in Table VI strongly validate the efficacy of our lifecycle module in bolstering the tracker robustness against mismatch scenarios (occlusion, etc.). The decrease in time consumption (-8.4ms) also proves the marked improvement in trajectory quality. As demonstrated in Table VII, our proposed score refinement exhibits a substantial enhancement in accuracy compared to the original confidence-based method [15], with a noteworthy AMOTA (+3.1%) increase. Besides, the proposed power prediction results in more precise score estimations, thereby contributing to improved accuracy (+0.4% AMOTA, +0.1% MOTA).
The effect of the Parallelization. Without any accuracy loss and without violating the online tracking principle, the multi-processing parallelization effectively alleviates the inherent serial logic of the TBD framework and improves the inference speed by 8.2% (-3.1ms).
@@ IV-E Hyperparameter Sensitivity Analysis
The plethora of hand-crafted hyperparameters poses a fundamental limitation to the TBD framework. In response, we conduct a sensitivity analysis on newly introduced hyperparameters on the nuScenes val set. Retaining other parameters, the linear search rule is executed for filter length $l_{lw}$, voxel mask size $\theta_{vm}$, decay rate $\sigma$ and match threshold $\theta_{fm}$. A critical finding in Fig. 5 is that Fast-Poly showcases insensitivity across all parameters barring $\theta_{fm}$, exhibiting an average fluctuation of less than 0.4% AMOTA. The pronounced decrease in $\theta_{vm}$ at 1 stems from erroneous matches induced by artificially narrowing the matching space. Furthermore, $\theta_{fm}$ emerges as a pivotal factor for all trackers, manifesting sensitivity across categories [1, 3, 17]. These findings underscore the resilience of our proposed method in parameter selection. The optimal hyperparameters are delineated in Section IV-A.

### v2

@@ III-D Parallelization
Although the TBD framework is inherently serial, certain modules can be parallelized. Consequently, in Fast-Poly, after tracking in the previous frame, motion estimation prediction and preprocessing are synchronized using multi-processing technology. Specifically, before the online tracking, two parallel processes $P_{1},P_{2}$ are initiated. This initialization also includes a shared flag $Flag$ and variable $V$ to synchronize status and transmit data between processes. At each frame $t$, $P_{1}$ filters raw detections while $P_{2}$ concurrently predicts the previous alive trajectory and implements online tracking. After completing their respective tasks, $P_{1}$ and the prediction part of $P_{2}$ enter a sleep state, remaining dormant until they are simultaneously awakened in the subsequent frame. Fast-Poly adheres to online tracking principles and fundamentally differs from [4], which employs multi-processing for synchronous multi-scene tracking.
@@ IV-A Dataset and Implementation Details
Metrics. We present three primary indicators: MOTA [21], AMOTA [10], and FPS (Frame Per Second), supplemented by three secondary indicators: IDS (ID Switch), FP, and FN. The latency of each experiment is the average time cost under multiple evaluations. We only measure the tracking time for all methods, excluding the detection time.
TABLE I: A comparison between our proposed method with other advanced methods on the nuScenes test set. This leaderboard is available at nuScenes official benchmark. $\ddagger$ means the GPU device. The reported runtimes of all methods exclude the detection time. Fast-Poly and [1] rely entirely on the detector input, as they do not utilize any visual or deep features during tracking.
[…]
@@
TABLE III: A comparison of the run-time performance on different devices on nuScenes val set with CenterPoint.
[…]
@@
[…]
@@
TABLE V: The ablation studies of each module on the nuScenes val. AG means A-GIoU. VM means Voxel Mask. LW means LightWeight Filter. CL means Confidence-Count Mixed Lifecycle. MP means Multi-Processing.
[…]
Baseline [1] 73.1 61.9 5.6 (188.7ms) 232 13051 17593
Exp1 73.1 62.4 19.9 (50.3ms) 342 13955 16794
Exp2 73.3 63.0 21.3 (46.9ms) 385 12732 17909
Exp3 73.3 62.9 21.7 (46.1ms) 427 14124 16523
Exp4 73.7 63.2 26.5 (37.7ms) 414 14713 15900
Exp5 73.7 63.2 28.9 (34.6ms) 414 14713 15900
TABLE VI: A comparison on distinct lifecycle modules on nuScenes val set. Average means using the online average score to delete. Latest means using the latest score to delete. Max-age means using the continuous mismatch time to delete. Other settings are under the best performance.
[…]
Count & Max-age 73.3 62.9 23.0 16523
Confidence [14] & Latest 70.6 63.2 45.8 19192
Confidence [14] & Average 73.3 63.1 28.3 16826
Confidence (Ours) & Average 73.7 63.2 28.9 15900
IV-B Run-time Discussion
The robustness of device migration. As demonstrated in Table III, we evaluated Fast-Poly across three CPU tiers. Since our method is learning-free, tracking accuracy remains unaffected by device migration. On the personal-level CPU, Fast-Poly exhibits the best real-time performance (29 FPS). Furthermore, without engineering acceleration, Fast-Poly can achieve nearly real-time performance (9 FPS) on limited CPU resources (Jetson Orin AGX), highlighting its practicality.
The parallel efficiency of Fast-Poly. As depicted in Fig. 1, when the number of objects increases significantly, the frame timing of Fast-Poly exhibits minimal increments ($\leq$30ms), highlighting the outstanding computational efficiency and feasibility of our method in autonomous driving scenarios.
TABLE VII: A comparison between our proposed method with other advanced methods on the nuScenes val set. $\ddagger$ means the GPU device. All metrics in the competition papers are reported. Methods in lines 1-7, lines 8-9, and line 10-13 use CenterPoint [19], LargeKernel3D [25] and DETR3D [37].
[…]
@@ IV-B Run-time Discussion
CC-3DT [38] 3090$\ddagger$ 35.9 32.6 9.6 2152
PF-Track [18] A100$\ddagger$ 36.2 300
Poly-MOT [1] 7945HX 38.7 32.3 16.3 926
Fast-Poly (Ours) 7945HX 40.7 33.1 39.3 1312
@@ IV-C Comparative Evaluation
On the val set, we employ CenterPoint [19] as the detector for a fair comparison. As shown in Table VII, Fast-Poly outperforms most learning-based/free methods by a significant margin in terms of both tracking accuracy (73.7% AMOTA, 64.2% MOTA) and latency (29 FPS). We surpass the baseline [1] with a notable improvement of +1.3% in MOTA and +0.6% in AMOTA accuracy, while 5$\times$ faster under identical settings. Fast-Poly exhibits stronger tracking performance with a powerful detector LargeKernel3D [25]. Utilizing the multi-camera detector DETR3D [37] with constrained performance, Fast-Poly shows robust real-time performance and accuracy, exceeding the end-to-end tracker PF-Track (+4.5% AMOTA) and TBD tracker CC-3DT [38] (+4.8% AMOTA). Notably, Poly-MOT requires a strict score filter threshold to achieve optimal AMOTA, reducing the latency advantage of our method.
Waymo. As shown in Table  IV, with the same detector, we exceed CasTrack [12] with a +1% improvement in MOTA on both sets, showcasing superior tracking performance. Without using the state-of-the-art detector, Fast-Poly ranks 4th on the Waymo tracking leaderboard among all LiDAR-only online methods with 63.6% MOTA to date. Despite the scarcity of methods showcasing real-time performance on Waymo, we believe that our inference speed (35.5 FPS) is competitive.
@@ IV-D Ablation Studies
The effect of the Proposed Metric A-GIoU. As demonstrated in Tables II and V, $A\text{-}gIoU$ results in a significant reduction in latency (-138.4ms), particularly during the overlap (-24.7ms) and convex hull (-83ms) calculations. $A\text{-}gIoU$ maintains AMOTA (73.1%) while improving MOTA (+0.5%). Fig. 3(a) shows this refinement stems from focusing on tight positional correlations, recalling more observations (-799 FN).
The effect of the Voxel Mask. As illustrated in Table V, the voxel mask improves overall accuracy (+0.2% AMOTA, +0.6% MOTA) and reduces latency (-3.4ms), verifying its effectiveness in avoiding invalid cost calculations. Specifically, it enhances absolute position constraints, preventing wrong matches from redundant detections (-1223 FP).
The effect of the Lightweight Filter. Table V indicates combining the lightweight filter enhances real-time performance (-0.8ms). However, the inner limited-step sliding window discards long-term temporal information for time-invariant states, slightly impacting estimation quality and overall tracking accuracy (-0.1% MOTA).
The effect of the Confidence-Count Mixed Lifecycle. Table V shows our lifecycle module (EXP4) increases processing speed by 4.8 FPS over EXP3, indicating more efficient tracklet management. It also improved AMOTA and MOTA by 0.4% and 0.3% respectively, validating its efficacy. As shown in Table VI, our tracklet termination strategy (line 3) outperforms the original score refinement [14] (line 2) with a 2.7% AMOTA increase and 2366 fewer FN, boosting tracker robustness against mismatch scenarios (occlusion, etc.). The smoother score prediction (line 4) further enhances tracking performance (+0.4% AMOTA, +0.1% MOTA, -926 FN).
The effect of the Parallelization. Without compromising accuracy or violating online tracking principles, the multi-processing parallelization effectively alleviates the inherent serial logic of the TBD framework and improves the inference speed by 8.2% (-3.1ms).
@@ IV-E Hyperparameter Sensitivity Analysis
The plethora of hand-crafted hyperparameters poses a fundamental limitation to the TBD framework. In response, we conduct a sensitivity analysis on newly introduced hyperparameters on the nuScenes val set. Retaining other parameters, the linear search rule is executed for filter length $l_{lw}$, voxel mask size $\theta_{vm}$, decay rate $\sigma$ and match threshold $\theta_{fm}$. A critical finding in Fig. 5 is that Fast-Poly showcases insensitivity to all parameters except $\theta_{fm}$, with an average fluctuation of less than 0.4% in AMOTA. The pronounced decrease in $\theta_{vm}$ at 1 stems from erroneous matches induced by artificially narrowing the matching space. Furthermore, $\theta_{fm}$ emerges as a pivotal factor for all trackers, manifesting sensitivity across categories [1, 3, 16]. These findings underscore the resilience of our proposed method in parameter selection. The optimal hyperparameters are delineated in Fig. 4.